a photon has a frequency of 8.63 x 10^8 Hz. What is the energy of this photon? (h=6.626 x10^-34 J X S)
1
Expert's answer
2020-04-03T11:49:51-0400

Solution.

"E＝h \times \nu"

"E＝8.63 \times 10^8 \times 6.626 \times 10^{-34}＝5.72 \times 10^{-25} \ J"

Answer.

"E＝5.72 \times 10^{-25} \ J"